An investor owns a security that is expected to return 10 percent in a booming economy and 3 percent in a normal economy. The overall expected return on the security is 5.45 percent. Given there are only two states of the economy, what is the probability that the economy will boom?

Multiple Choice

  • 28 percent

  • 45 percent

  • 33 percent

  • 41 percent

  • 35 percent
Answer #1
Assuming that probability of boom economy is x therefore probability of normal economy = 1-x
expected return % = 5.45%
Therefore we have below equation -
x* 10% + (1-x) * 3% = 5.45%
10%x+3%-3%x = 5.45%
7%X = 2.4500%
X= 35.00%
therefore correct answer = 35%